A R1 000 000 trust fund was set up and to be used by an 8-year old nephew when he goes to college. In 8 years, how much will the fund be if the investment rate is 7.5% compounded quarterly?

1
Expert's answer
2021-11-23T17:50:33-0500

The amount earned on the principal amount P

P invested for t

t years at the rate of r

r compounded quarterly is:

"A=P(1+\frac{r}{4})^{4t}"

Now, we are given:

"P=1000000\\t=8\\r=7.5\%=\frac{7.5}{100}=0.075"

Putting the values of "P=1000000 , t=8, r=0.075"

"A=P(1+\frac{r}{4})^{4t}"

we get:

"=1000000(1+\frac{0.075}{4})^{4×8}\\=1000000(1.01875)^{32}\\=1812023.793"


Hence, the investment will be Php 1812023.793 at the end of 8 years.